in a computer class of 42 students, 28 like to work on the hardware, 24 like to work on the software while 6 neither like to work on the hardware or software.
How many like to work on both hardware and software?
How many like to work on hardware only?
How many like to work on software only?

Answer by CubeyThePenguin(3113)

Draw a Venn Diagram:

hardware | both | software | neither
-------------------------------------------------
28 ______| x___ | 24 ________ | 6

(28 - x) + x + (24-x) + 6 = 42
52 - x = 36

x = 16

like to work on both = x = 16
work on hard ware = 28 - x = 12
work on soft ware = 24 - x = 8